OpenStudy (anonymous):

Outcome Payout HHH $10 HHT $3 HTH $0 HTT -$3 TTT $10 THH -$3 THT $0 TTH $0 Kara is playing a game where she flips a coin 3 times. She wins and loses different amounts of money based on the outcomes. What is the EXPECTED VALUE for the game?

OpenStudy (anonymous):

Expected value is the sum of the probability of each even times the outcome of the event. So she wins $10 with probability 2/8, $3 with probability 1/8, $0 with probability 3/8, and loses $3 with probability 2/8. So, E = 10(2/8) + 3(1/8) + 0(3/8) - 3(2/8) =5/4 + 3/8 - 3/4 =7/4 = $1.75

OpenStudy (anonymous):

Crap!

OpenStudy (anonymous):

here you mean?

OpenStudy (anonymous):

I see it...10(2/8) = 10/4

OpenStudy (anonymous):

yeah...stupid arithmetic error

OpenStudy (anonymous):

each of these has probability \[\frac{1}{8}\] so just compute \[10+3-3+10-3\] and divide by 8

OpenStudy (anonymous):

i get \[\frac{17}{8}\] where did the 4 come from?

OpenStudy (anonymous):

Ack! Too early. That'll teach me not to stay up so late playing video games.

OpenStudy (anonymous):

in a situation like this think of "expected value" as the average. add up and divide by total

OpenStudy (anonymous):

Yeah, i should have. But I've conditioned myself to do it the full way since I've had sadistic teachers who throw in unfair coins into the mix.

OpenStudy (anonymous):

but these are the only answer choices A) $0 B) $2.13 C) $3.29 D) $4.14